xXMarcelieXx:

I dont get this.. how are they all with the same degree 3?

xXMarcelieXx:

1 attachment
xXMarcelieXx:

@Zepdrix

Nnesha:

\[\large\rm x^3y'+8x^2yy'+4xy^2-y^3=0\] ignore y'\[\large\rm x^\color{Red}{3} +8x^\color{Red}{2}y^\color{Red}{1} +4x^\color{Red}{1}y^\color{Red}{2}-y^\color{red}{3}=0\] \[\rm first~term \rightarrow x^\color{red}{3} , ~degree ~3\]\[\rm 2nd~term \rightarrow 8x^\color{red}{2}y^\color{red}{1} \rightarrow 2+1=3, ~~ degree~ 3\]\[\rm 3rd~term \rightarrow 4x^\color{Red}{1}y^\color{Red}{2} \rightarrow 1+2=3 , degree 3\]\[\rm last~term \rightarrow y^\color{red}{3} , degree 3\]